5. Suppose X ~ Poisson(A = 5) and Y ~ Poisson(λ = 10), and they are independent. Using the moment generating function method, find the distribution of Z = X + Y.
Answer #1

here as we know that mgf of poisson distribution: M(t)=e^{lambda (e^{t}-1) }

therefore mgf of X: Mx(t)=5(et-1)

mgf of Y: Mx(t)=10(e-1)

therefore mgf of Z =Mz(t)=MX+Y(t)=MX(t)*MY(t) =5(et-1)*10(e-1) =15(e-1)

for mgf of Z follows Poisson distribution mgf with parameter lambda =15

hence distribution of Z is Poisson with parameter lambda =15